OpenStudy (anonymous):

Determine the number and type of solutions for x^2 + 5x + 6 = 0.

OpenStudy (anonymous):

\[x^2+5x+6=0\] a=1 b=5 c=6

OpenStudy (mertsj):

\[b^2-4ac=25-4(1)(6)=25=24=1\]

OpenStudy (mertsj):

Discriminant is positive. How many times does it touch the x axis?

OpenStudy (mertsj):

Whoops. 25-24

OpenStudy (anonymous):

So yes, with a positive discriminant there would be two "real" solutions. :)

OpenStudy (campbell_st):

since the value of the descrimant >0 and a square number... there are two distinct rational roots